OpenStudy (anonymous):

solve the equation: 3-n/2=3/4

OpenStudy (anonymous):

\[\frac{6-n}{2} = \frac{3}{4}\] Multiply 2 both the sides and let us see what you get...

OpenStudy (anonymous):

@nique504 If you are only here to look here and there and you don't wanna reply then you will not get the solution..

OpenStudy (anonymous):

multiply both sides by 2?

OpenStudy (anonymous):

That I have already told you.. After multiplying what you get??

OpenStudy (anonymous):

nevermind, i don't need help solving this.

OpenStudy (anonymous):

Then you can close the question.. Nevermind..

OpenStudy (anonymous):

shut up.haha